OpenStudy (anonymous):

7 2/3 - 1 3/5 = 5 1/8

jhonyy9 (jhonyy9):

23/3 - 8/5 = 41/8 23*40-24*8=15*41 920-172=748 not equal 15*40=600+15=615 - so result that this not is equal